A rectangular room is 1.9 times as lomng as it is wide, and its perimeter is 33 meters. Find the dimension of the room. The length is what in meters? And the width is what in meters?
Answer by checkley77(12844)

L=1.9W
P=2L+2W
33=2(1.9W)+2W
33=3.8W+2W
5.8W=33
W=33/5.8
W=5.69 METERS FOR THE WIDTH.
L=1.9*5.69=10.81 METERS FOR THE LENGTH.
PROOF:
33=2*10.81+2*5.69
33=21.62+11.38
33=33